OpenStudy (anonymous):

A person weighing 785 newtons on the surface of the earth would weigh 298 newtons on the surface of Mars. What is the magnitude of the gravitational field strength on the surface of Mars?

OpenStudy (awesome781):

sounds like a physics question..

OpenStudy (anonymous):

yh

OpenStudy (fifciol):

\[W_1=mg_{earth}\]\[W_2=mg_{mars}\] Since the mass cannot change anywhere( unless you lost or gain your weight): \[\frac{ W_1 }{ g_{earth} }=\frac{W_2}{ g_{mars} } \Rightarrow g_{mars}=\frac{ W_2 }{ W_1}g_{earth}\] w1=785, w2=298, g earth=9.91

OpenStudy (anonymous):

Thanks that was so helpful.

OpenStudy (fifciol):

yw:)